3y-(4-2y)=3(y+2)
Can u please help me with this its not in a text book its a worksheet

Answer by Lacey020991(49)

3y-(4-2y)=3(y+2)
---Your goal is to get the y all by itself.
---First you would distribute everything
3y -4 +2y = 3y +6
---Next you would combine like terms
5y -4 = 3y +6
---Next you would subtract 3y from each side. On the right, the +3y and the -3y would cancel leaving you with 6.
2y -4 = 6
---Then you would add 4 to each side. Again, the -4 and +4 would cancel leaving you with 2y
2y = 10
---Then you would divide by 2 on each side.
y=5
Always check your answer!
3*5 - (4-2*5) = 3(5+2)
15 - (-6) = 3(7)
21 = 21 check!